# Linear viscoelastic rheological FrBD models

### v1

@@ Abstract
In [1], a new modeling paradigm for developing rate-and-state-dependent, control-oriented friction models was introduced. The framework, termed Friction with Bristle Dynamics (FrBD), combines nonlinear analytical expressions for the friction coefficient with constitutive equations for bristle-like elements. Within the FrBD framework, this letter introduces two novel formulations based on the two most general linear viscoelastic models for solids: the Generalized Maxwell (GM) and Generalized Kelvin-Voigt (GKV) elements. Both are analyzed in terms of boundedness and passivity, revealing that these properties are satisfied for any physically meaningful parametrization. An application of passivity for control design is also illustrated, considering an example from robotics. The findings of this letter systematically integrate rate-and-state dynamic friction models with linear viscoelasticity.
@@ 1 Introduction
A prominent example is LuGre [6, 7], originally introduced in the context of control-oriented modeling to account for pre-sliding displacement and frictional lag through an internal bristle state. Since its introduction, the LuGre model and its variants have been widely adopted and extended in both theoretical and applied studies, particularly in control design and compensation schemes [2, 3, 4, 5, 11]. Despite its success, the LuGre model exhibits well-known limitations, most notably the lack of passivity for general parameter choices and operating conditions, unless additional velocity-dependent damping terms are introduced [12, 13, 14]. This has prompted ongoing research into alternative formulations that preserve physical consistency whilst remaining amenable to analysis and control synthesis [8, 9, 10].
@@
Within this context, the present letter extends the FrBD framework by introducing and analyzing a class of nonlinear viscoelastic models inspired by classical linear rheological elements. In particular, formulations based on the Generalized Maxwell (GM) and Generalized Kelvin-Voigt (GKV) representations are developed, allowing for richer descriptions of friction dynamics in solids. The mathematical properties of the resulting model are rigorously investigated, with particular emphasis on existence and uniqueness of solutions, boundedness, and passivity. Beyond theoretical analysis, the letter also examines the qualitative dynamical behavior of the proposed models, showing that they reproduce key friction phenomena such as pre-sliding displacement and frictional hysteresis observed in experimental studies. Finally, the practical relevance of the approach is demonstrated through a control application involving a robotic manipulator, where passivity of the friction model is sapiently exploited for output-feedback control design. This letter develops a unified framework combining general linear viscoelastic theory with rate-and-state-dependent friction laws, resulting in a family of dynamic models that admit clear physical interpretation and systematic characterization using standard engineering methods [22].
@@
Finally, an application of Theorem 1.1 below will guide the derivation of the rate-and-state-dependent friction models developed in the letter.
@@ 2 Model derivation
In turn, approximating [^1] The approximation is needed to deal with the non-differentiability of the friction force at zero. Informally, for a regularized model, it may be justified by observing that the derivative of $f_{\textnormal{r}}(v)$ at zero will be very large, and can be replaced by $\lim_{v\to 0}\mu(v)/\mathinner{\!\left\lvert v\right\rvert}_{\varepsilon}=\mu(0)/\varepsilon$. [1, 23]
@@ 3.2 Mathematical properties
This section investigates the boundedness and passivity properties of the FrBDn+1 models (12) and (13). Formal mathematical definitions of these notions are well established in the literature (see, e.g., [18]) and are therefore not repeated here. Instead, their physical interpretation in the context of friction modeling is briefly discussed. In particular, boundedness reflects the physically reasonable expectation that the friction force and internal states remain finite, coherently with (7). On the other hand, passivity captures the intrinsically dissipative nature of friction, and should be satisfied by any bona fide friction model [13]. Essentially, it states that the power exchanged by the friction force must be nonnegative, that is, $pf(t)v(t)\geq 0$, where $p\in\mathbb{R}_{>0}$ is the normal force. For the FrBDn+1 models presented in this paper, both boundedness and passivity are always guaranteed.
@@
$\begin{split}\dot{V}(t)&=pf(t)v(t)-\dfrac{p\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}}{\mu\bigl(v(t)\bigr)}f^{2}(t)\\ &\quad-\sum_{i=1}^{n}\dfrac{p}{\bar{c}_{i}}(\bar{k}_{i}z_{i}(t)-\bar{k}_{0}z_{0}(t))^{2},\quad t\in(0,T).\end{split}$ (21)
The above inequalities (18) and (21) will be instrumental to prove boundedness and passivity for the FrBDn+1 models (12) and (13). The main results in this sense are enounced in the following Sections 3.2.1 and 3.2.2.
@@ 3.2.1 Boundedness
$\begin{split}\dot{V}(t)&=\gamma_{1}\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\sqrt{V(t)}\\ &\quad-p\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\Biggl(\dfrac{f^{2}(t)}{\varrho_{2}}+\sum_{i=1}^{n}\dfrac{p}{\varrho_{1}\tau_{i}\bar{k}_{i}}f_{i}^{2}(t)\Biggr)\\ &\leq-\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\sqrt{V(t)}\Bigl(\gamma_{2}\sqrt{V(t)}-\gamma_{1}\Bigr),\quad t\in(0,T).\end{split}$ (23)
@@ 4 Dynamical behavior and control applications
The dynamics of the FrBDn+1 models is investigated qualitatively in Section 4.1 for $n=1$. Indeed, for $n=1$, both the GM and GKV variants collapse into a Standard Linear Solid element, for which analytical interconversions are explicitly available in the literature [21]. A control application related to a robotic manipulator is then exemplified in Section 4.2.
@@ 4.1 Dynamical behavior
The dynamical behavior of the FrBD2 models is discussed concerning pre-sliding displacement and frictional lag
4.1.1 Pre-sliding displacement
Prior to breakaway, microscopic relative motion occurs as the reactive friction force builds up. This initial regime is commonly referred to as the pre-sliding displacement. Courtney-Pratt and Eisner showed that, when the applied external force remains below the breakaway threshold, friction exhibits a behavior analogous to that of a nonlinear spring [19].
To assess whether the FrBD model reproduces this characteristic, a first set of simulations was performed by postulating the friction coefficient as
$\mu(v_{\textnormal{s}})=\mu_{\textnormal{d}}+(\mu_{\textnormal{s}}-\mu_{\textnormal{d}})\mathrm{e}^{-(\mathinner{\!\left\lvert v_{\textnormal{s}}\right\rvert}/v_{\textnormal{S}})^{\delta}},$ (25)
where $\mu_{\textnormal{d}},\mu_{\textnormal{s}}\in\mathbb{R}_{>0}$ denote the dynamic and static friction coefficient, respectively, $v_{\textnormal{S}}\in\mathbb{R}_{\geq 0}$ is the Stribeck velocity, $\delta\in\mathbb{R}_{\geq 0}$ the Stribeck exponent. In these simulations, a unit-mass system was subjected to a sinusoidal external force with an amplitude equal to 90% of the breakaway force, whilst varying the excitation frequency over $\omega=1$, 5, and 10 Hz. The resulting friction force $pf$ is plotted as a function of displacement $x$ in Figure 3(a) . The simulated response shows good qualitative agreement with the experimental observations of Courtney-Pratt and Eisner [19], and closely matches the behavior predicted by the LuGre model [7]. Figure 3 was generated using the model parameters listed in Table 1, with $\sigma_{0}=\bar{k}_{1}$, $\sigma_{1}=\tau_{1}(\bar{k}_{0}+\bar{k}_{1})$, and $\gamma_{1}=\tau_{1}$ in (5) according to the GM parametrization.
@@ 4.1.1 Pre-sliding displacement
$\delta$ Stribeck exponent - 2
@@
4.1.2 Frictional lag
Hess and Soom [20] investigated the velocity-dependent dynamic behavior of friction during unidirectional motion. Their experiments demonstrated the presence of hysteresis in the friction force-velocity relationship: for a given velocity magnitude, the friction force is lower during deceleration than during acceleration. In addition, the width of the hysteresis loop was observed to vary with the rate of change of velocity, expanding or contracting as this rate increases.
For relatively high values of the damping coefficient $\sigma_{1}$, the hysteresis loop is generally inverted when the Stribeck velocity $v_{\textnormal{S}}$ is sufficiently large. However, the conventional hysteretic behavior is recovered for sufficiently small values of $v_{\textnormal{S}}$. Figure 3(b) shows the hysteresis curves obtained by imposing relative velocities at different excitation frequencies, $\omega=25$, 50, and 100 Hz. As the frequency increases, the peak friction force $pf$ decreases, whilst the hysteresis loops become progressively broader, consistent with the preceding discussion.
Figure 3: Dynamical behavior of the FrBD2 models: (a) Pre-sliding displacement hysteresis; (b) frictional lag hysteresis
@@ 4.2 Feedback control of a robotic arm
Combining (30) and (31) implies $s\in L^{2}(\mathbb{R}_{>0};\mathbb{R})$. Invoking similar arguments as in [11], it may be finally concluded that $\lim_{t\to\infty}\tilde{q}(t)=0$. The result is formalized in Theorem 4.1 below.
@@
Consider the ODE system (26) with the observer (28). Then, the control input (27) ensures that $\lim_{t\to\infty}\tilde{q}(t)=0$.
@@ 5 Conclusion
Utilizing a natural energy-based Lyapunov interpretation, the proposed models, formulated as systems of $n+1$ nonlinear ODEs, were shown to be well-posed and to satisfy boundedness and passivity for all parametrizations. Qualitative analysis demonstrated that key friction phenomena, including pre-sliding displacement and frictional hysteresis, are accurately reproduced and consistent with experimental observations and established dynamic friction models. Finally, a robotic control example illustrated how the passivity of the FrBD model can be effectively exploited in output-feedback control design.
These findings confirm the FrBD framework as a physically grounded and control-oriented approach to dynamic friction modeling. In particular, the results of this letter unify, for the first time, linear viscoelasticity and rate-and-state friction models. Future extensions may incorporate nonlinear rheological representations.

### v2

@@ Abstract
Recently, a new modeling paradigm, termed Friction with Bristle Dynamics (FrBD), has been introduced for developing rate-and-state-dependent, control-oriented friction models. The framework combines nonlinear analytical expressions for the friction coefficient with constitutive equations for bristle-like elements. Within the FrBD framework, this letter introduces two novel formulations based on the two most general linear viscoelastic models for solids: the Generalized Maxwell (GM) and Generalized Kelvin-Voigt (GKV) elements. Both are analyzed in terms of boundedness and passivity, revealing that these properties are satisfied for any physically meaningful parametrization. An application of passivity for control design is also illustrated, considering an example from robotics. The findings of this letter systematically integrate rate-and-state dynamic friction models with linear viscoelasticity.
@@ 1 Introduction
A prominent example is LuGre [6, 7], originally introduced in the context of control-oriented modeling to account for pre-sliding displacement and frictional lag through an internal bristle state. Since its introduction, the LuGre model and its variants have been widely adopted and extended in both theoretical and applied studies, particularly in control design and compensation schemes [2, 3, 4, 5, 11]. Despite its success, the LuGre model exhibits well-known limitations, most notably the lack of passivity for general parameter choices and operating conditions, unless velocity-dependent damping terms are introduced [12, 13, 14]. Additionally, the LuGre model involves several parameters whose physical interpretation is not straightforward, which complicates its use in multibody simulations [23]. This has prompted ongoing research into alternative formulations that preserve physical consistency whilst remaining amenable to analysis and control synthesis [8, 9, 10].
@@
Within this context, the present letter extends the FrBD framework by introducing and analyzing a class of nonlinear viscoelastic models inspired by classical linear rheological elements. In particular, formulations based on the Generalized Maxwell (GM) and Generalized Kelvin-Voigt (GKV) representations are developed, allowing for richer descriptions of friction dynamics in solids. The mathematical properties of the resulting model are rigorously investigated, with particular emphasis on existence and uniqueness of solutions, boundedness, and passivity. Beyond theoretical analysis, the letter also examines the qualitative dynamical behavior of the proposed models, showing that they reproduce key friction phenomena such as frictional hysteresis and relaxation behaviors observed in experimental studies. Finally, the practical relevance of the approach is demonstrated through a control application involving a robotic manipulator, where passivity of the friction model is sapiently exploited for output-feedback control design. This letter develops a unified framework combining general linear viscoelastic theory with rate-and-state-dependent friction laws, resulting in a family of dynamic models that admit clear physical interpretation and systematic characterization using standard engineering methods [21]. The proposed formulations are particularly relevant for applications involving highly viscoelastic materials, including polymers and rubbers.
@@
Finally, an application of Theorem 1.1 below will guide the derivation of the friction models developed in the letter.
@@ 2 Model derivation
In turn, approximating [1, 22]
@@ 3.2 Mathematical properties
This section investigates the boundedness and passivity properties of the FrBDn+1 models (12) and (13). Formal mathematical definitions of these notions are well established in the literature (see, e.g., [18]) and are therefore not repeated here. Instead, their physical interpretation in the context of friction modeling is briefly discussed. In particular, boundedness reflects the physically reasonable expectation that the friction force and internal states remain finite, coherently with (7). On the other hand, passivity captures the intrinsically dissipative nature of friction, and should be satisfied by any bona fide friction model [13]. Essentially, it states that the power exchanged by the friction force must be nonnegative, that is, $pf(t)v(t)\geq 0$, where $p\in\mathbb{R}_{>0}$ is the normal component of the contact force. For the FrBDn+1 models presented in this paper, both boundedness and passivity are always guaranteed.
@@
$\begin{split}\dot{V}(t)&=pf(t)v(t)-\dfrac{p\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}}{\mu\bigl(v(t)\bigr)}f^{2}(t)\\ &\quad-\sum_{i=1}^{n}\dfrac{p}{\bar{c}_{i}}\bigl(\bar{k}_{i}z_{i}(t)-\bar{k}_{0}z_{0}(t)\bigr)^{2},\quad t\in(0,T).\end{split}$ (21)
The above (18) and (21) will be instrumental to prove boundedness and passivity for the FrBDn+1 models (12) and (13). The main results in this sense are enounced in the following Sections 3.2.1 and 3.2.2.
@@ 3.2.1 Boundedness
$\begin{split}\dot{V}(t)&=\gamma_{1}\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\sqrt{V(t)}\\ &\quad-p\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\Biggl(\dfrac{f^{2}(t)}{\varrho_{2}}+\sum_{i=1}^{n}\dfrac{1}{\varrho_{1}\tau_{i}\bar{k}_{i}}f_{i}^{2}(t)\Biggr)\\ &\leq-\mathinner{\!\left\lvert v(t)\right\rvert}_{\varepsilon}\sqrt{V(t)}\Bigl(\gamma_{2}\sqrt{V(t)}-\gamma_{1}\Bigr),\quad t\in(0,T).\end{split}$ (23)
@@ 4 Dynamical behavior and control applications
The dynamics of the FrBDn+1 models is investigated qualitatively in Section 4.1, whereas a control application related to a robotic manipulator is then exemplified in Section 4.2.
@@ 4.1 Dynamical behavior
The dynamical behavior of the FrBD models is discussed concerning frictional lag and relaxation.
4.1.1 Frictional lag
Hess and Soom [19] investigated the velocity-dependent dynamic behavior of friction during unidirectional motion. Their experiments demonstrated the presence of hysteresis in the friction force-velocity relationship: for a given velocity magnitude, the friction force is lower during deceleration than during acceleration. In addition, the width of the hysteresis loop was observed to vary with the rate of change of velocity, expanding or contracting as this rate increases. To assess whether the FrBD model reproduces this characteristic, a first set of simulations was performed by postulating the friction coefficient as
$\mu(v_{\textnormal{s}})=\mu_{\textnormal{d}}+(\mu_{\textnormal{s}}-\mu_{\textnormal{d}})\mathrm{e}^{-(\mathinner{\!\left\lvert v_{\textnormal{s}}\right\rvert}/v_{\textnormal{S}})^{\delta_{\textnormal{S}}}},$ (25)
where $\mu_{\textnormal{d}},\mu_{\textnormal{s}}\in\mathbb{R}_{>0}$ denote the dynamic and static friction coefficient, respectively, $v_{\textnormal{S}}\in\mathbb{R}_{\geq 0}$ is the Stribeck velocity, $\delta_{\textnormal{S}}\in\mathbb{R}_{\geq 0}$ the Stribeck exponent.
For relatively high values of the damping coefficient $\sigma_{1}$, the hysteresis loop is generally inverted when the Stribeck velocity $v_{\textnormal{S}}$ is sufficiently large. However, the conventional hysteretic behavior is recovered for sufficiently small values of $v_{\textnormal{S}}$. Figure 3 shows the hysteresis curves obtained by imposing relative velocities at different excitation frequencies, $\omega=25$, 50, and 100 Hz. As the frequency increases, the peak friction force $pf$ decreases, whilst the hysteresis loops become progressively broader, consistent with the preceding discussion. Figure 3 was generated using the model parameters listed in Table 1, with $\sigma_{0}=\bar{k}_{1}$, $\sigma_{1}=\tau_{1}(\bar{k}_{0}+\bar{k}_{1})$, and $\gamma_{1}=\tau_{1}$ in (5) according to the GM parametrization.
Figure 3: Frictional lag hysteresis.
@@ 4.1.1 Frictional lag
$\delta_{\textnormal{S}}$ Stribeck exponent - 2
@@
4.1.2 Relaxation
When a constant relative sliding velocity is imposed between a viscoelastic body and a rigid substrate, the generated friction force does not instantaneously attain its steady-state value. Instead, a transient phase is observed before the stationary friction level is fully established. This phenomenon is commonly referred to as relaxation behavior. For highly viscoelastic materials, such as the polymers typically used in tyres, internal stress relaxation occurs over a broad frequency spectrum. Accurately capturing this behavior requires rheological models that incorporate multiple relaxation times [20]. Figure 4 illustrates the relaxation behavior predicted by the FrBDn+1-GM models for $n=1$, 2, and 3, with parameters as in Table 1, $\tau_{1}=1$, $\tau_{2}=0.1$, $\tau_{3}=0.01$ s and $v=0.01$ $\textnormal{m}\,\textnormal{s}^{-1}$. According to (14) and (12), all the variants yield the same steady-state response, but their relaxation dynamics is influenced by the rheological order.
Figure 4: Relaxation behavior for different rheological orders.
@@ 4.2 Feedback control of a robotic arm
Combining (30) and (31) implies $s\in L^{2}(\mathbb{R}_{>0};\mathbb{R})$. Invoking similar arguments as in [11], it may be finally concluded that $\lim_{t\to\infty}\tilde{q}(t)=0$. The claim, formalized in Theorem 4.1 below, is the logical consequence of known results on passive systems interconnections [24].
@@
Consider the ODE system (26) with the observer (28). Then, the control input (27) ensures that $\lim_{t\to\infty}\tilde{q}(t)=0$ for all $(q_{0},\dot{q}_{0},z_{0},f_{1,0},\dots,f_{n,0})\in\mathbb{R}^{n+3}$.
@@ 5 Conclusion
Utilizing a natural energy-based Lyapunov interpretation, the proposed models, formulated as systems of $n+1$ nonlinear ODEs, were shown to be well-posed and to satisfy boundedness and passivity for all parametrizations. Qualitative analysis demonstrated that key friction and rheological phenomena, including frictional hysteresis and stress relaxation, are accurately reproduced and consistent with experimental observations and established dynamic models. Finally, a robotic control example illustrated how the passivity of the FrBD model can be effectively exploited in output-feedback control design. These findings confirm the FrBD framework as a physically grounded and control-oriented approach to dynamic friction modeling. In particular, the results of this letter unify, for the first time, linear viscoelasticity and rate-and-state friction models. Future research will address experimental validation and explore nonlinear rheological extensions and integration within multibody simulations.